OpenStudy (anonymous):

The soccer team is designing a triangular flag for their games. The length of the base is 5 inches more than the height, h. Express the area of the flag as a function of the height. Answer choices: A) A(h)=h^2+5/2h B) A(h)=1/2h^2-5/2h C) A(h)=1/2h^2+5/2 D) A(h)=1/2h^2+5/2h Please help anyone :)

OpenStudy (anonymous):

b = h + 5 Area of triangle = bh/2 = (h + 5)h / 2 = (h^2 + 5) / 2 = 1/2h^2 + 5/2 I believe it's C?